The longer side of a rectangle is three times the length of the shorter side. If the length of the diagonal is 10 cm, find the dimensions of the rectangle.
Answer by checkley77(12844)

X^2+3X^2=10^2
4X^2=100
X^2=100/4
X^2=25
X=SQRT25
X=5 ANS.
PROOF:
5^2+3^5^2=10^2
25+3*25=100
25+75=100
100=100